OpenStudy (anonymous):

Rewrite the following expression as the tangent of an angle. tan23 + tan55/1-(tan23)(tan55)

OpenStudy (anonymous):

Please clarify your expression. We don't know if you mean tan(23) + (tan(55))/(1-tan(23)tan(55)) or something else.

OpenStudy (anonymous):

The sum formula, tan (x+y) = tan(x) + tan (y) all divided by 1 - tan (x) tan(y) So your expression is the formula for tan (x+y) where x = 23 and y = 55 So, final answer; tan (78)

OpenStudy (anonymous):

Your welcome.

OpenStudy (jdoe0001):

http://home.windstream.net/okrebs/Ch10-6.gif

OpenStudy (anonymous):

Thank you. I wasn't sure, so I had to ask!

OpenStudy (anonymous):

No problem.